OpenStudy (anonymous):

find the square root of 625? show work

OpenStudy (anonymous):

\[\sqrt{625} =\sqrt{25*25}=25\]

OpenStudy (amistre64):

they said find it; not write it :) we have to delve into proofs and stuff maybe